if i travel 40 miles an hour and go 8 miles , how much time will it take me tio get to work?
Answer by Maths68(1474)

Distance= 8 miles
Speed = 40mph
Time = ?
Time = Distance/speed
Time = 8/40
Time = 1/5 hour
Time = 1/5 * 60
Time = 60/5 * 60
Time = 12 minutes